### v1

@@ Abstract
This study investigates the timing and spectral characteristics of X-ray bursts from the neutron star system EXO 0748-676 using the NuSTAR observatory’s FPMA and FPMB instruments in the 3-79 keV range. We identify Type I X-ray bursts driven by thermonuclear explosions on the neutron star’s surface, notably a significant burst at , indicating rapid energy release, followed by a recoil burst at , reflecting stabilization. The correlation between burst timing and the neutron star’s optical period suggests modulation by its rotation and periodic accretion dynamics. Spectral modeling reveals a photon index of and a cutoff energy of , indicating a hot corona around the neutron star. The measured flux of approximately underscores the dynamic nature of accretion-driven systems. Calculated luminosities derived from distance estimates range from to . Comparative analysis with prior observations from the IBIS/ISGRI instrument on the INTEGRAL satellite shows variability in emission characteristics, including softer photon indices and higher cutoff energies in 2003 and 2004. Our examination of smaller energy gaps (3-7 keV, 7-12 keV, etc.) reveals energy-dependent behavior in burst characteristics, enhancing our understanding of nuclear burning phases. Overall, these findings validate models describing Type I X-ray bursts and lay the groundwork for future investigations into similar astrophysical systems and stellar evolution processes in extreme environments.
Keywords NuSTAR accretion accretion discs-stars neutron-pulsars individual EXO 0748-676 X-rays low mass x-ray binaries (LMXBs)
@@ 3.1 Timing analysis
After completing the above steps, the light curve was plotted by selecting the relevant time intervals and applying binning to improve the time resolution, using the lcurve command. The light curve was then divided into different energy bands to analyze energy-dependent variations in the count rate. Energy filters corresponding to specific ranges (e.g., 3–7 keV, 7–12 keV, etc.) were applied to split the light curve into segments, allowing for a detailed study of the source’s behavior across different energy ranges and detecting burst events. This was done using the nuproducts command, where specific energy ranges were defined in terms of PI channels, followed by the same steps as above using lcmath for background subtraction and barycorr for barycentric correction for all the energy bins. In Fig. 2, the time curve for the full energy range (3–79 keV) is shown alongside smaller energy ranges (3–7 keV, 7–12 keV, up to 30–79 keV) to examine the nature of the outburst in EXO 0748-676, enabling a detailed analysis of how the outburst evolves across different energy bands.
@@
This analysis of the X-ray burst data in the 3-79 energy range, combined with the neutron star’s optical period of 13,750 seconds Parmar et al. (1986); Knight et al. (2022); Southwell et al. (1996), reveals key aspects of the system’s behavior. It supports the interpretation that the burst represents a significant thermonuclear event, while the smaller bursts, such as the one at , likely result from less energetic fluctuations related to the star’s accretion system or magnetospheric activity. These findings align with the well-documented behaviors of Type I X-ray bursts, where accretion, nuclear burning, and rotational dynamics all interplay to generate recurrent, explosive energy releases.
@@
The intervals between these bursts provide further insights into the system’s dynamics. The gap between the first major burst at and the recoil burst at is approximately 984 seconds, reflecting rapid energy release and adjustment. More intriguingly, the gap between the burst at and the next significant burst at is roughly 13,778 seconds, closely matching the neutron star’s optical period of 13,750 seconds. This suggests a potential link to the star’s rotational cycle, where some bursts could be influenced by periodic accretion events associated with its binary interaction, pointing to the role of rotational dynamics and the magnetic field structure in modulating the burst mechanism.
@@
The energy range over which X-ray bursts are observed plays a significant role in understanding the physical mechanisms driving these thermonuclear explosions on neutron stars. In the broad 3-79 keV energy range, the burst at stands out as the most energetic event, with a count rate of 1615.88 counts per second. This range captures the full evolution of the burst, from the initial thermonuclear ignition to the cooling phase, providing a complete picture of the burst dynamics. The broad energy spectrum shows not only the intense peak but also the subsequent smaller bursts, such as the recoil burst at , with a count rate of 32.14. These secondary bursts likely result from residual energy dissipation or magnetospheric adjustments. However, when we break down the burst behavior into narrower energy ranges, important patterns emerge. In the soft X-ray band (3-7 keV), the primary burst shows a much lower count rate of 647.47 counts per second, indicating that much of the burst’s energy is emitted in higher energy bands. Soft X-rays primarily capture the cooling phase of the burst, which decays more slowly, highlighting the prolonged emission from the neutron star’s surface. This behavior aligns with our understanding of Type I X-ray bursts, where initial high-energy photons rapidly decline, leaving softer X-rays to dominate the emission for longer periods Galloway et al. (2008b).
@@
In contrast, mid-energy ranges (7-12 keV and 12-18 keV) show bursts with intermediate intensity, emphasizing the importance of these bands in capturing the peak of the thermonuclear explosion. For instance, in the 7-12 keV range, the primary burst reaches 752.16 counts per second, while in the 12-18 keV range, it drops to 189.09 counts per second. This suggests that the bulk of the burst’s energy is emitted in these mid-energy bands, where nuclear burning on the neutron star’s surface releases a significant amount of X-rays before cooling. The physics here is driven by helium burning, which ignites under conditions of high pressure and temperature on the star’s surface, leading to rapid energy release followed by radiative cooling Strohmayer and Bildsten (2003); Fowler and Hoyle (1965); Lewin et al. (1997, 1993).
@@
Higher energy bands (18-30 keV and 30-79 keV) capture the hardest X-rays, but the intensity is much lower compared to the full energy range. For instance, in the 18-30 keV band, the primary burst has a count rate of 36.57 counts per second, while in the 30-79 keV band, it further decreases to 6 counts per second. These hard X-rays are likely emitted during the early phases of the burst, reflecting the hottest parts of the thermonuclear explosion. The rapid decline in higher energy bands supports the idea that hard X-rays are primarily linked to the initial stages of nuclear burning, with less emission as the system cools. Such behavior is consistent with the high-energy photons emitted during the peak of a photospheric radius expansion (PRE) burst, where radiation pressure lifts the outer layers of the neutron star before settling into a cooling phase Strohmayer (2006); Strohmayer and Bildsten (2003).
@@ 3.2 Phase-averaged Spectral analysis
Spectral parameters | NuSTAR(2024) | IBIS | |
ISGRI(2003) | ISGRI(2004) | ||
6.350.15 | - | - | |
Photon Index | 1.240.014 | 1.30.4 | 1.60.4 |
Cutoff Energy (keV) | 36.201.04 | 4415 | 5015 |
1.77E-024.39E-04 | - | - | |
1864.80 | - | - | |
Degrees of freedom | 1899 | - | - |
0.98 | - | - | |
Flux | 381.170.016 | 320150 | 421150 |
The observed spectrum was fitted using a model combination of CONSTANT PHABS CUTOFFPL, which includes three key components. The PHABS model accounts for photoelectric absorption due to neutral interstellar matter, characterized by the hydrogen column density (). The CUTOFFPL model is a power-law with an exponential cutoff, typically used to describe the X-ray emission in accreting systems, especially those containing compact objects like neutron stars. The spectral fit resulted in reduced chi-squared values () of 0.98 for FPMA and 1.007 for FPMB, indicating that the model accurately describes the data, as illustrated in Fig. 3.
The fit revealed several key parameters. The hydrogen column density was found to be , suggesting significant absorption along the line of sight, likely due to dense interstellar matter or local absorbing material. The photon index () indicates that the X-ray spectrum is relatively hard, pointing to non-thermal processes or Comptonized thermal photons in a hot corona around the neutron star. The cutoff energy () reflects the energy at which the X-ray emission begins to decline sharply, possibly due to thermal Comptonization processes in the corona. The flux in the 3-79 keV energy range was measured at , providing a useful reference for future observations.
These results were compared with previous observations by the IBIS/ISGRI instrument on the INTEGRAL satellite, which observed the source in the 20-100 keV range in 2003 and 2004. The earlier IBIS/ISGRI data showed slightly softer photon indices, with values of (2003) and (2004), suggesting the emission was somewhat softer during those observations. Additionally, the cutoff energies were higher, 44 keV in 2003 and 50 keV in 2004, compared to NuSTAR’s 36.2 keV, indicating a potentially hotter Comptonizing region or corona in those periods. The flux values also varied between in 2003 and in 2004, reflecting typical fluctuations in accretion-driven systems. The ISGRI detector, designed for hard X-ray to soft gamma-ray observations, is well-suited for studying sources like EXO 0748-676 during both outburst and quiescent phases. Their spectral fitting was performed using XSPEC v11.2, and a 10% systematic error was added to account for uncertainties in their spectral extraction and response matrix. The fitting was done using a cutoff power-law model, where the photon index () and cutoff energy () parameters matched the general trends seen in NuSTAR’s observations.
@@
Based on the gravitational redshift, Özel suggested that the distance of EXO 0748-676 is 9.2 1.0 kpc Özel (2006). From a strong X-ray burst, Wolff et al. Wolff et al. (2005) derived a distance of 7.7 kpc for a helium burst. Galloway et al. analyzed several type I X-ray bursts from EXO 0748-676 and estimated a distance of 7.4 kpc Galloway et al. (2008b), different from the value of 9.2 kpc reported by Özel Özel (2006). Taking into account the touchdown flux and the high binary inclination of EXO 0748-676, Galloway, Özel, and Psaltis recently gave a distance of 7.1 1.2 kpc Galloway et al. (2008a); Zhang et al. (2011).
Using a calculated flux of , the corresponding luminosities for these distances were derived. For a distance of 9.2 kpc, the luminosity is . At a distance of 7.7 kpc, the luminosity is . For a distance of 7.4 kpc, the luminosity is , with a negligible error due to the exact distance value. Finally, for a distance of 7.1 kpc, the calculated luminosity is . These results demonstrate the relationship between distance and luminosity, with greater distances resulting in higher luminosity values.
The luminosities derived from different distance estimates of EXO 0748-676 in our study can be compared to those reported in Trigo et al. (2006); Boirin et al. (2007). Using the data from revolution 719, Díaz Trigo et al. Trigo et al. (2006) report an absorbed flux in the range of 0.6–10 keV measured at . This absorbed flux, when corrected for no attenuation, corresponds to an unabsorbed flux of . Boirin et al. Boirin et al. (2007) further calculated an unabsorbed bolometric flux of across the 0.1-100 keV energy band. At a distance of 5 kpc, this flux corresponds to a luminosity of , while at 10 kpc, it yields a significantly higher luminosity of .
In contrast, our analysis using the calculated flux of erg/cm²/s provides a range of luminosities based on various distance estimates. For a distance of 9.2 kpc, the luminosity is erg/sec, and at 7.7 kpc, it is erg/sec. As the distance decreases to 7.4 kpc, we find a luminosity of erg/sec, and at 7.1 kpc, the luminosity is erg/sec.

### v2

@@ Abstract
This study investigates the timing and spectral characteristics of X-ray bursts from the neutron star system EXO 0748-676 using the NuSTAR observatory’s FPMA and FPMB instruments in the 3-79 keV range. A notable type I X-ray thermonuclear burst at , with a peak of 1,615.88 counts per second, was observed, revealing key characteristics of helium ignition on the neutron star’s surface. Distinct energy dependencies were found, with soft X-ray emissions (3-7 keV) associated with the cooling phase and mid-energy bands (7-18 keV) capturing the bulk of the energy during ignition. Hard X-rays (18-79 keV) were linked to the early stages of the explosion. The study also examined the timing of bursts, uncovering correlations between burst intervals and the neutron star’s optical period, suggesting a link between accretion processes and the system’s binary nature. Analysis of the eclipse profile revealed energy-dependent ingress and egress durations, providing insights into the density gradient of the surrounding material. By applying two spectral continuum models, the research illuminated the complex emission mechanisms, with one model highlighting Comptonization and the other emphasizing both thermal and non-thermal components. Comparisons with historical data suggest temporal changes in the accretion environment, reflected in variations in photon indices and cutoff energies. The calculated flux in a full energy range of 3-79 KeV using both the continuum models are found to be . Calculated luminosities derived from different distance estimates range from to . The findings contribute to a deeper understanding of accretion processes, thermonuclear dynamics, and the behavior of neutron stars, with implications for future investigations into the complex interactions in low-mass X-ray binaries and high-energy astrophysical phenomena.
Keywords NuSTAR accretion accretion discs-stars neutron-pulsars individual EXO 0748-676 X-rays low mass x-ray binaries (LMXBs) Eclipse
@@ 3.1 Timing analysis
After completing the previous steps, the light curve was plotted by selecting the relevant time intervals and applying binning to improve time resolution, using the lcurve command. Figure 2 shows the light curve of the LMXB pulsar EXO 0748-676 in the energy range of 3–79 keV. The plot uses a logarithmic scale for the count rate axis, allowing clear visualization of the eclipse and the prominent burst events in the source EXO 0748-676. In Fig. 2, the red lines represent the most significant burst observed at (MJD 20478.519486587) in the neutron star EXO 0748-676 during NuSTAR observation ID 91001324002. The same burst is also shown magnified in the top-right panel to highlight its detailed structure. Clearly, from Fig. 2, five distinct eclipse cycles can be observed, the properties of which will be discussed in more detail in a later section.
The light curve was then further divided into different energy bands to analyze energy-dependent variations in the count rate. Energy filters corresponding to specific ranges (e.g., 3–7 keV, 7–12 keV, etc.) were applied to split the light curve into segments, allowing for a detailed study of the source’s behavior across different energy ranges and detecting burst events. This was done using the nuproducts command, where specific energy ranges were defined in terms of PI channels, followed by the same steps as above using lcmath for background subtraction and barycorr for barycentric correction for all the energy bins. In Fig. 3, the time curve for the full energy range (3–79 keV) is shown alongside smaller energy ranges (3–7 keV, 7–12 keV, and up to 30–79 keV) to examine the nature of the outburst in EXO 0748-676. This enables a detailed analysis of how the outburst evolves across different energy bands. Note that the actual height of the significant burst is clipped to make smaller outbursts more visible.
@@
This analysis of the X-ray burst data in the 3-79 energy range, combined with the neutron star’s optical period of 13,767.5 seconds , reveals key aspects of the system’s behavior. It supports the interpretation that the burst represents a significant thermonuclear event, while the smaller bursts, such as the one at , likely result from less energetic fluctuations related to the star’s accretion system or magnetospheric activity. These findings align with the well-documented behaviors of Type I X-ray bursts, where accretion, nuclear burning, and rotational dynamics all interplay to generate recurrent, explosive energy releases.
@@
The intervals between these bursts provide further insights into the system’s dynamics. The gap between the first major burst at and the recoil burst at is approximately 984 seconds, reflecting rapid energy release and adjustment. More intriguingly, the gap between the burst at and the next significant burst at is roughly 13,778 seconds, closely matching the neutron star’s optical period of 13,767.5 seconds. This suggests a potential link to the star’s rotational cycle, where some bursts could be influenced by periodic accretion events associated with its binary interaction, pointing to the role of rotational dynamics and the magnetic field structure in modulating the burst mechanism.
@@
Fig. 4 shows the significant Type I X-ray burst observed at MJD 20478.519486587 during the NuSTAR observation, displayed across different energy ranges. The energy range over which X-ray bursts are observed plays a significant role in understanding the physical mechanisms driving these thermonuclear explosions on neutron stars. In the broad 3-79 keV energy range, the burst at stands out as the most energetic event, with a count rate of 1615.88 counts per second. This range captures the full evolution of the burst, from the initial thermonuclear ignition to the cooling phase, providing a complete picture of the burst dynamics. The broad energy spectrum shows not only the intense peak but also the subsequent smaller bursts, such as the recoil burst at , with a count rate of 32.14. These secondary bursts likely result from residual energy dissipation or magnetospheric adjustments. However, when we break down the burst behavior into narrower energy ranges, important patterns emerge. From Fig. 4, in the soft X-ray band (3-7 keV), the primary burst shows a much lower count rate of 687.71 counts per second, indicating that much of the burst’s energy is emitted in higher energy bands. Soft X-rays primarily capture the cooling phase of the burst, which decays more slowly, highlighting the prolonged emission from the neutron star’s surface. This behavior aligns with our understanding of Type I X-ray bursts, where initial high-energy photons rapidly decline, leaving softer X-rays to dominate the emission for longer periods Galloway et al. (2008b).
@@
In contrast, mid-energy ranges (7-12 keV and 12-18 keV) show bursts with intermediate intensity, emphasizing the importance of these bands in capturing the peak of the thermonuclear explosion. For instance, in the 7-12 keV range, the primary burst reaches 799.14 counts per second, while in the 12-18 keV range, it drops to 211.78 counts per second. This suggests that the bulk of the burst’s energy is emitted in these mid-energy bands, where nuclear burning on the neutron star’s surface releases a significant amount of X-rays before cooling. The physics here is driven by helium burning, which ignites under conditions of high pressure and temperature on the star’s surface, leading to rapid energy release followed by radiative cooling Strohmayer and Bildsten (2003); Fowler and Hoyle (1965); Lewin et al. (1997, 1993).
@@
Higher energy bands (18-30 keV and 30-79 keV) capture the hardest X-rays, but the intensity is much lower than the full energy range. For instance, in the 18-30 keV band, the primary burst has a count rate of 43.28318 counts per second, while in the 30-79 keV band, it further decreases to 1.87 counts per second. These hard X-rays are likely emitted during the early phases of the burst, reflecting the hottest parts of the thermonuclear explosion. The variation in the amplitude of the Type I X-ray burst count rate with energy is plotted in Fig. 5. The rapid decline in higher energy bands supports the idea that hard X-rays are primarily linked to the initial stages of nuclear burning, with less emission as the system cools. Such behavior is consistent with the high-energy photons emitted during the peak of a photospheric radius expansion (PRE) burst, where radiation pressure lifts the outer layers of the neutron star before settling into a cooling phase Strohmayer (2006); Strohmayer and Bildsten (2003).
Eclipse profile
From Fig. 2, five distinct eclipse cycles were observed in the NuSTAR observation of the LMXB EXO 0748-676. Each eclipse cycle, caused by the low-mass main-sequence secondary star UY Vol passing in front of the neutron star, has a period of seconds, or approximately 3.824 hours, which closely matches the previously reported orbital period of EXO 0748-676 at 3.824 hours Wade et al. (1985); Knight et al. (2022); Southwell et al. (1996). The eclipse typically lasts about seconds, equivalent to minutes. On average, the eclipse ingress/egress time is found to be about 12 seconds. To understand more about the eclipse and its energy dependency, we have folded the extracted light curve about the orbital period of 3.824 hours and normalized it throughout the light curve with the average count rate of the burst-free and out-of-eclipse light curve so that the eclipse profile now has an out-of-eclipse count rate of 1.0, and the eclipse totality leveled at 0.0. The eclipse profile is extracted for different energy ranges and is plotted in Fig. 6. Note that the eclipse profile at different energy ranges is set to a vertical offset for visual clarity. From Fig. 6, it is observed that initially during ingress, there is a gradual decrease in the count rate towards totality, and the same phenomenon is observed at egress, where the count rate gradually rises to the normalized count rate.
For a more detailed analysis, we further divided the ingress and egress phases into two-time frames, defined as and . Here, and represent the points where the count rate reaches 10% and 90% of the average burst-free, out-of-eclipse count rate, respectively. For the ingress phase, marks the onset of the count rate’s gradual decline, while denotes its approach to the out-of-eclipse count rate level.
To account for random fluctuations, these times are defined as the points at which the average count rate first surpasses the specified percentage and remains there for at least five seconds Knight et al. (2022). The measured values of and across different energy ranges, from 3–7 keV to 30–79 keV, are displayed in the upper panels of Fig. 7, showing ingress on the left and egress on the right, with the corresponding ingress/egress durations shown in the lower panels.
As seen in Fig. 7, the ingress start time increases for higher energy ranges from 3–7 keV to 12–18 keV, decreases for the 18–30 keV range, and then rises again in the 30–79 keV range. Similarly, the ingress end time shows a slight decrease from 3–7 keV to 7–12 keV, followed by a linear increase with higher energy ranges. This pattern indicates that ingress begins later at higher photon energies but ends quickly, whereas the lower energy ranges show longer ingress durations, particularly within the 12–18 keV range, with durations decreasing towards the 3–7 keV region.
The observed behavior of the ingress start time and end time across different energy ranges in EXO 0748-676 reflects how the neutron star’s environment and radiation characteristics vary with photon energy. As seen in Fig. 7, ingress Start Time decreases for higher energy ranges from 3–7 keV to 12–18 keV, increases for the 18–30 keV range, and then again decreases for the 30–79 keV range. The decrease in at higher photon energies, from 3–7 keV up to 12–18 keV, suggests that the high-energy X-rays encounter the obscuring material later in the ingress phase. Higher energy photons can penetrate the neutron star’s surrounding environment more effectively, making them visible longer as the eclipse starts. For the 18–30 keV range, the increase in may indicate a transitional energy regime where the photons interact more strongly with the denser regions near the neutron star, becoming obscured sooner. Similarly, the subsequent decrease in for the 30–79 keV range might indicate that high-energy X-rays, although affected by scattering or absorption processes, still penetrate more effectively than intermediate-energy X-rays. These high-energy photons are likely interacting with less dense material, allowing them to be visible longer at the start of the ingress.
Similarly, the ingress end time shows a slight increase from 3–7 keV to 7–12 keV, followed by a linear decrease with higher energy ranges. The initial increase in from 3–7 keV to 7–12 keV suggests that lower-energy photons are more susceptible to absorption by material in the eclipse, causing the count rate to drop faster and earlier in the process. And the linear decrease in with increasing photon energy indicates that high-energy photons end the ingress phase more quickly. This is because high-energy photons can traverse the surrounding material more effectively, which results in a sharper and more rapid end to the ingress phase for higher energy bands.
The pattern suggests that the ingress duration varies with energy. At lower energies, X-rays are more likely absorbed or scattered by the material around the neutron star, resulting in a longer ingress period at the energy range of 3-7 KeV to 12-18 KeV. At higher energies, 18 to 79 KeV, photons are less affected by scattering and absorption, which shortens the ingress duration. The shorter ingress duration at high energies implies that these photons can more directly reflect the neutron star’s emissions and are obscured only by the densest parts of the surrounding material.
The egress phase shows an interesting variation across energy ranges, reflecting changes in the timing of when photons of different energies begin and end their exit from the eclipse. The egress start time () slightly decreases from 3–7 keV to 7–12 keV, indicating that the eclipse ends earlier for these slightly higher energies. This trend then reverses as begins to increase linearly with energy, suggesting that for higher photon energies, the egress starts progressively later. This delay at higher energies could imply that these photons are initially blocked by denser material and only gradually become visible as the line of sight clears. Meanwhile, the egress end time () generally decreases from 3–7 keV to 12–18 keV, meaning the eclipse duration shortens with energy in this range, causing higher energy photons to be visible sooner. However, then shows a slight increase in the 18–30 keV range and becomes relatively constant at higher energies (30–79 keV), possibly due to higher energy photons passing through remaining obstructions more effectively.
Examining the overall duration of egress, it tends to decrease with increasing energy, suggesting that higher-energy photons are less affected by residual obstructions as the eclipse ends. A slight increase in egress duration is observed in the 18–30 keV range, indicating more complex interactions between these mid-energy photons and the material surrounding the source. Overall, this pattern reveals that higher-energy photons experience a shorter egress duration, likely because they penetrate through the material more effectively as it clears from the line of sight. This energy-dependent behavior of the ingress/egress phase offers insights into the density distribution and composition of the material surrounding the LMXB, highlighting how various photon energies respond to different extents of absorption and scattering.
In the 3-79 keV band, the observed ingress and egress durations for EXO 0748-676 are 24.89 and 27.99 seconds, respectively. These durations reflect the time taken for the neutron star to enter and exit an eclipse, providing insight into the material surrounding the companion. Comparing this with a study conducted in the 0.2-10 keV range for the XMM-Newton observation of EXO 0748-676, it appears that eclipse behavior is energy-dependent, suggesting that an absorbing region surrounds the companion star and may be undergoing ablation. At lower energies, photons interact more with the absorbing material, while in higher-energy bands, such as 3-79 keV, the duration is shorter due to better photon penetration. This previous study also inferred a high neutron star mass and hard equation of state (EoS), suggesting the system’s compactness and gravitational effects influence these eclipse timings. Such comparisons across energy ranges allow for a more detailed understanding of the system’s density, inclination, and mass properties Knight et al. (2022).
In our analysis of the total eclipse duration, characterized by and , we observe distinct trends across various energy ranges, offering valuable insights into the dynamics of the eclipse. The total duration is calculated as , where represents either 10 or 90. Figure 8 illustrates the variation in total eclipse duration across different energy ranges. From Fig. 8, for , we observe a decrease from the energy range of 3-7 keV to 7-12 keV, suggesting that at these lower energy levels, the eclipse duration is shorter, indicating a rapid transition as the neutron star obscures the X-ray source. Beyond 12 keV, gradually and linearly increases with energy. This could be due to increased complexity in interactions at higher photon energies, potentially related to variations in the density or structure of the material surrounding the neutron star.
In contrast, exhibits a decrease from the 3-7 keV range to the 18-30 keV range, indicating that the onset of the gradual decline in count rate occurs earlier in these energy bands. This rapid decline in the observable count rate could be attributed to higher photon absorption or scattering processes occurring in the material surrounding the neutron star. After the 18-30 keV range, begins to increase, suggesting that the eclipse duration becomes prolonged at higher energies. This could be due to increased photon penetration through the absorbing material or the effects of relativistic beaming, where high-energy photons are emitted in a more focused manner.
Furthermore, the difference gradually decreases with increasing energy, indicating that as photon energy rises, the duration difference between the initial onset of the count rate decline and its approach to the out-of-eclipse level becomes smaller. This observation may suggest that higher-energy photons experience a more uniform interaction with the surrounding material, leading to a more consistent eclipse duration. These trends in the duration of the eclipse phases, influenced by energy levels, provide valuable information about the structure and behavior of the material surrounding the neutron star, including its density and composition.
@@ 3.2 Phase-averaged Spectral analysis
Parameter | Component | Unit | Value | Confidence Range (90%) |
Constant | - | (frozen) | - | |
Phabs | ||||
CutoffPL | - | |||
CutoffPL | keV | |||
CutoffPL |
This study explores EXO 0748-676’s emission characteristics through two distinct spectral continuum models applied to data obtained from the NuSTAR telescope: the first continuum model (Model 1), which uses a constant * phabs * cutoffpl configuration, and the second continuum model (Model 2), a more complex model that incorporates constant * phabs * highecut (powerlaw + bbodyrad). Each model reveals key physical insights into the binary system’s structure and provides a comparative perspective on the mechanisms that govern its emission.
Model 1 uses a cutoff power-law to describe the spectrum, along with interstellar absorption modeled by phabs. This model represents the X-ray emission as a power-law with an exponential cutoff, suitable for characterizing Comptonized emission from an optically thin, high-temperature electron plasma located near the neutron star. The spectral fit resulted in reduced chi-squared values () of 0.982 for FPMA and 1.006 for FPMB, indicating that the model accurately describes the data, as illustrated in Fig. 9. Best fitted Spectral Parameters for EXO 0748-676 with 90% Confidence Errors are listed in Table 2. The fitted value for the hydrogen column density () is , consistent with previous observations and indicating substantial absorption, likely due to both interstellar matter and circumstellar material around the binary. The power-law photon index () of suggests a hard spectrum, indicative of high-energy emissions from the electron plasma. Additionally, the high-energy cutoff parameter () is found to be , suggesting that most of the energy is emitted below this threshold, which points to the temperature of the corona surrounding the neutron star. Finally, the power-law normalization () of defines the relative flux of this component. Together, these parameters provide a reliable depiction of the system’s X-ray continuum emission, primarily shaped by the high-energy scattering processes occurring in the hot corona.
Parameter | Component | Unit | Value | Confidence Range (90%) |
Constant | - | (frozen) | - | |
Phabs | ||||
HighEcut | keV | |||
HighEcut | keV | |||
Powerlaw | - | |||
Powerlaw | ||||
BBodyRad | keV | |||
BBodyRad | - |
Model 2 offers a more sophisticated interpretation, incorporating both non-thermal and thermal emission components, which gives a comprehensive view of the system. This model includes a combination of a power-law, similar to that in Model 1, with an additional blackbody radiation component representing thermal emission from the neutron star’s surface or the inner accretion disk. Here, the cutoff and folding energies are modeled using the highecut component, which captures the high-energy break more flexibly. The spectral fitting yielded reduced chi-squared values of for FPMA and for FPMB, suggesting that the model effectively represents the data, as shown in Fig. 10. The optimal spectral parameters for EXO 0748-676, along with their 90% confidence errors, are detailed in Table 3.
The hydrogen column density () in this model is , which is slightly lower than in Model 1. The model’s cutoff energy () is , and the folding energy () is , parameters that characterize the Comptonized corona emission and indicate possible variations in the temperature and density of the electron cloud surrounding the neutron star. The power-law photon index () of suggests a particularly hard spectrum, indicative of strong Compton scattering effects within a highly energetic environment.
A notable feature of Model 2 is the inclusion of a blackbody radiation component. This component, with a temperature () of , reflects thermal emission, possibly originating from the neutron star’s heated surface due to accretion activity. This thermal component represents a blackbody surface temperature on the order of , which provides insight into the neutron star’s thermal processes and how accretion impacts the surface. The power-law normalization () is , and the blackbody normalization () is , showing the respective contributions of each emission type within the spectrum. This model, with its thermal component, gives a more nuanced view of the spectral properties, separating out the potential emission sources, including the neutron star surface, the boundary layer, and the surrounding accretion disk.
The two models provide complementary insights into the behavior of EXO 0748-676. Model 1 employs a single-component approach that effectively captures the continuum emission, suggesting that the system’s high-energy behavior is predominantly influenced by scattering processes within a hot, optically thin corona. In contrast, Model 2 offers a more nuanced exploration of the spectral composition by isolating both thermal and non-thermal contributions, indicating the likely presence of a corona alongside a heated neutron star surface. This two-component model points to a multi-temperature structure within the binary system, with the lower cutoff energy in Model 2 potentially reflecting cooler regions within the accretion environment or variations in accretion dynamics over time. Overall, both models yield valuable insights, with Model 1 providing a robust representation of continuum emission, while Model 2 enhances our understanding of the thermal and non-thermal processes involved. The ability of Model 2 to distinguish between these components is particularly relevant for elucidating the physical conditions surrounding the neutron star, including the interplay between accretion disk dynamics, coronal heating, and surface emission. Such spectral analysis is crucial for probing the physical conditions governing EXO 0748-676, illuminating its high-energy emissions, and contributing to a broader understanding of low-mass X-ray binary systems and neutron star characteristics. Further observations across different energy bands and over extended periods would be beneficial for refining these models and enhancing our comprehension of the spectral complexity and temporal variability inherent to EXO 0748-676.
We have also calculated the flux in a full energy range of 3-79 KeV using both the continuum models with 90% confidence errors. The first model provided a flux of . This model suggests a significant contribution from non-thermal processes, indicating the influence of high-energy emissions in the observed spectrum. Conversely, the second model yielded a slightly lower flux of . The minor difference in flux between the two models highlights the variability in the spectral shape and the contributions of thermal and non-thermal emissions. Overall, both models underscore the complexity of the accretion processes and emission mechanisms present in EXO 0748-676, providing crucial insights into the neutron star’s environment and the underlying physics governing low-mass X-ray binaries.
The spectral fits obtained from NuSTAR observations were compared with earlier results from the IBIS/ISGRI instrument on the INTEGRAL satellite, which observed EXO 0748-676 in the 20-100 keV range in 2003 and 2004 Götz et al. (2006). These earlier observations showed slightly softer photon indices compared to NuSTAR’s, with values of in 2003 and in 2004. This indicates that the emission was somewhat softer during those periods. Softer photon indices typically suggest that the spectrum contains more lower-energy photons, implying differences in the energy distribution of the emitted radiation between the observations. In contrast, NuSTAR’s observations revealed a harder spectrum, suggesting a more compact and energetic emission region.
In addition to the differences in photon indices, the cutoff energies observed by INTEGRAL in 2003 and 2004 were 44 keV and 50 keV, respectively. These values are higher than NuSTAR’s observed cutoff energy of 36.2 keV, which suggests that the Comptonizing region or corona was cooler during the NuSTAR observations. A lower cutoff energy means that the high-energy photons are less frequent, which could point to a less energetic particle population responsible for upscattering the photons. The differences in cutoff energies could be linked to variations in the accretion environment or changes in the physical conditions around the neutron star at different times.
The flux values observed by IBIS/ISGRI varied over time, with measurements of in 2003 and in 2004. These values reflect the typical fluctuations seen in accretion-driven systems, where changes in the accretion rate can lead to variations in the observed flux. In comparison to the 2004 data, the flux values obtained from NuSTAR’s spectral fits are lower. For Model 1, the flux was , while for Model 2, the flux was slightly lower at . These flux values indicate that EXO 0748-676 was dimmer during the NuSTAR observations compared to the earlier measurements from 2004. The decrease in flux may suggest reduced accretion activity or changes in the system’s viewing angle.
@@
To further understand the physical properties of EXO 0748-676, the distances to the source were considered. Several distance estimates have been proposed in the literature. Özel et al. suggested a distance of 9.2 1.0 kpc based on the gravitational redshift of the system Özel (2006). Wolff et al. derived a distance of 7.7 kpc from observations of a helium burst Wolff et al. (2005), while Galloway et al. analyzed several type I X-ray bursts and estimated a distance of 7.4 kpc Galloway et al. (2008b). More recently, Galloway, Özel, and Psaltis proposed a distance of 7.1 1.2 kpc based on the touchdown flux and the high binary inclination of the system Galloway et al. (2008a); Zhang et al. (2011). These varying distance estimates impact the calculation of the source’s luminosity, which depends on both the flux and the distance.
Distance (kpc) | Model 1 Luminosity (ergs/s) | Model 2 Luminosity (ergs/s) |
9.2 ± 1 | ||
7.7 ± 0.9 | ||
7.4 | ||
7.1 ± 1.2 |
Using the flux values obtained from both models, the luminosity of EXO 0748-676 was calculated for each of the proposed distances of 9.2 kpc, 7.7 kpc, 7.4 kpc, and 7.1 kpc. The calculated luminosity are detailed in Table 4. These results demonstrate a clear relationship between the distance and the luminosity of EXO 0748-676, with larger distances corresponding to higher luminosities. This relationship is a direct consequence of the inverse square law, where luminosity increases with the square of the distance for a given flux. The difference between the two models is relatively small, with Model 1 consistently yielding slightly higher luminosity values than Model 2. However, the uncertainty in the flux from Model 2 is larger, leading to broader error margins in the derived luminosities.
When comparing these results to previous luminosity estimates, we find consistency in the overall trends, though there are some notable differences. Díaz Trigo et al. Trigo et al. (2006) reported an absorbed flux of for EXO 0748-676 in the 0.6–10 keV range, when corrected for no attenuation, which corresponds to an unabsorbed flux of . This flux value is lower than the fluxes obtained from both NuSTAR models. Boirin et al. Boirin et al. (2007) further calculated an unabsorbed bolometric flux of across the 0.1-100 keV energy band which is greater than that predicted at NuSTAR. However, when considering a distance of 5 kpc, their data gives a luminosity of , which is comparable to the luminosities derived from both NuSTAR models at 7.4 kpc.